OpenStudy (anonymous):

27 ducks are marching to the pond. 5 lose their way, 13 return and 9 of them make it to the pond. What happens to the rest of them?

OpenStudy (anonymous):

many things may happen really so it depends

OpenStudy (anonymous):

Hint: How many are "the rest of them"?

OpenStudy (anonymous):

flew away, hunted

OpenStudy (shubhamsrg):

none of them could rest, all of them were moving :P

OpenStudy (anonymous):

d(^.^)b

OpenStudy (anonymous):

5 + 13 + 9 = 27... Something happened to all of the 27. So... nothing happened to the nonexistent "rest of them".

OpenStudy (shubhamsrg):

hmm..

OpenStudy (anonymous):

.-.

OpenStudy (shubhamsrg):

:D

OpenStudy (anonymous):

I don't think there is a joke lol.

OpenStudy (anonymous):

Yet I'm "lol"ing

OpenStudy (anonymous):

O.O

OpenStudy (anonymous):

i find this amusing :D

OpenStudy (shubhamsrg):

20 ? :|

OpenStudy (anonymous):

b + 2r = 62 b + r = 36 b + b = ? -------- Substitution: b + 2r = 62 -2r b = 62 - 2r (62 - 2r) + r = 36 62 - 2r + r = 36 62 - r = 36 -r = -26 r = 26 b + r = 36 b + (26) = 36 b = 10 Sooo b + b 10 + 10 20

OpenStudy (anonymous):

I even showed my work (;

OpenStudy (shubhamsrg):

(Y) :D

OpenStudy (anonymous):

We're awesome lol

OpenStudy (anonymous):

hmmm interesting cool

OpenStudy (anonymous):

gtg ty gn

OpenStudy (anonymous):

Bye @iforgot hope you enjoyed your stay lol ^.^

OpenStudy (anonymous):

yep loved it :D